# When Simultaneous Localization and Mapping Meets Wireless Communications: A Survey

### v1

@@ III-C Passive SLAM for Networked Autonomous Systems
Building on these pillars, the perception module of a networked autonomous system incorporates and extends the vision-based inference mechanisms that have shaped robotic SLAM. Networked implementations reinterpret them within a distributed framework, where visual cues are fused with radio-based signals and network state information to maintain situational awareness under dynamic conditions.
@@ III-D Active SLAM for Networked Autonomous Systems
Active SLAM is deemed a computationally intensive task, hence a natural question is when one should terminate active SLAM and switch to passive in order to allocate the computational resources to other tasks. Balancing active SLAM decisions and exogenous tasks is crucial, since in most real-world tasks, active SLAM is only a ‘flow duct‘to achieve an intended goal. Additionally, having stopping criteria is imperative, because it can be shown that after reaching a particular point, more information would lead not only to a diminishing return effect but also, in case of ambiguous or contradictory information, to an unrecoverable state (e.g., numerous incorrect loop closures). Lastly, another critical research direction concerns the establishment of mathematical guarantees for active SLAM and the derivation of near-optimal control policies. Given that obtaining an exact, closed form solution to the problem is in the general case intractable, developing approximation algorithms with well-defined performance bounds becomes essential. A representative example of this line of work is the application of submodularity theory [18] in the related domain of active sensor placement, which enables provable near-optimal solutions under certain conditions.
@@ IV-A Mathematical and Algorithmic Methods
Mathematical modeling plays an important role in bridging SLAM with wireless communications. In conventional SLAM, probabilistic estimation techniques help robots and autonomous vehicles navigate and reconstruct their environment. However, when applied to wireless environments, SLAM must account for additional challenges posed by signal propagation, multipath effects, and dynamic radio environments. This interplay between localization, mapping, and wireless signal processing requires advanced mathematical frameworks that incorporate elements of RF-based sensor fusion, probabilistic inference, signal estimation, and information vectorization.
@@ IV-A1 Probabilistic Approaches in SLAM
Extending previous studies on 5G SLAM, [7] explored localization techniques suitable for new radio (NR) and sub-THz 5G communications. They proposed an L1-regularized least-squares approach with low computational complexity to obtain sparse range-angle charts at individual sensing locations. In addition, a novel state model was introduced for tracking diffuse and specular scatterers using an interacting multiple model extended Kalman filter (IMM-EKF) and a smoother. The IMM design included a finite set of motion models with probabilistic transitions between them, following the principles of mmWave SLAM. Experimental results demonstrated the effectiveness of the proposed methods, with IMM-EKF-based tracking significantly outperforming static modeling approaches.
@@ IV-A3 Vector-field SLAM
A vector field is defined as a regular grid with fixed node positions $\mathbf{b}_{i}$, where $i=\{1,\dots,N\}$, while each node $\mathbf{m}_{i}\in\mathbb{R}^{M}$ represents a vector of expected signal values when the robot is positioned at $\mathbf{b}_{i}$ with a fixed orientation $\theta_{0}$. Without loss of generality, let $\theta_{0}=0$. The vector field and robot pose are then estimated throughout the SLAM challenge as follows.
@@
$\mathbf{x}_{t}=\begin{bmatrix}x(t)\\ y(t)\\ \theta(t)\end{bmatrix}^{T}.$
@@
$\mathbf{x}_{0}=\begin{bmatrix}0\\ 0\\ 0\end{bmatrix}^{T}.$
At each time step $t=\{1,\dots,T\}$, the robot executes a motion control input $\mathbf{u}_{t}$ with covariance $\mathbf{R}_{t}$ and receives a measurement $\mathbf{z}_{t}$ of the continuous robot state signals with covariance $\mathbf{Q}_{t}$.
@@
$\mathbf{x}_{t}=\mathbf{g}(\mathbf{x}_{t-1},\mathbf{u}_{t})+\mathbf{e}_{u},$
@@
$\mathbf{z}_{t}=\mathbf{h}\left(\mathbf{x}_{t},\begin{bmatrix}\mathbf{m}_{1},&\dots,&\mathbf{m}_{N}\end{bmatrix}^{T}\right)+\mathbf{e}_{z},$
@@
$h\left(\mathbf{x}_{t},\begin{bmatrix}\mathbf{m}_{1},&\dots,&\mathbf{m}_{N}\end{bmatrix}^{T}\right)$ $=$ $h_{R}\Bigg(h_{0}\Big(x,y,$
$\ \ \begin{bmatrix}\mathbf{m}_{1},&\dots,&\mathbf{m}_{N}\end{bmatrix}^{T}\Big),\theta\Bigg),$
@@
$h_{0}\left(x,y,\begin{bmatrix}\mathbf{m}_{1},&\dots,&\mathbf{m}_{N}\end{bmatrix}^{T}\right)=\sum_{j=0}^{3}w_{j}\mathbf{m}_{i_{j}}.$
@@
IV-B SLAM Techniques
@@ IV-B2 LiDAR-based SLAM
In [11], the authors investigated hardware and software co-design strategies to meet the requirements of mobile robot localization and object perception. To enable accurate mobile positioning, they proposed a SLAM strategy that integrates LiDAR, a monocular camera, and an IMU. For object perception, the study exemplifies the extraction of meter information using DL techniques for meter detection, localization, and text recognition. The integration of 5G and Wi-Fi communication enables an intelligent robotic inspection system with a network management platform. The authors of [129] addressed the challenge of robust long-term localization in cluttered indoor environments, such as offices and corridors, using LiDAR point clouds within an area graph. This hierarchical, topometric semantic map representation uses polygons to define spaces such as rooms, corridors, or buildings. Their approach, called area graph localization (AGloc), is augmented by a clutter removal subsampling method to remove excessive noise, as well as sampling techniques and weighting functions for global robot localization (Fig. 5). First, a set of candidate robot positions is generated using a sampling technique, after which a scoring function evaluates each candidate based on its match with the area graph. The point-to-point iterative closest point (ICP) algorithm and a weighting function enable localization using only wall and passage representations, ensuring robust lifelong localization despite the presence of non-static obstacles such as furniture, pets, and humans. Additionally, a ’corridorness’ score function is introduced to provide downsampled point clouds, improving ICP localization in corridor-like environments.
@@
In [152], a high-precision multimodal approach for indoor positioning was proposed using LiDAR as the central sensor modality. This method improves localization accuracy by incorporating widely available technologies, such as Wi-Fi access points (APs) and low-cost cameras. In this hybrid scheme, positioning data from LiDAR-based SLAM is used as a theoretical reference for visual images, while WiFi signals estimate the rough location through likelihood. Finally, RGB images refine the ground map and improve the localization accuracy. As mentioned above, LiDAR-based localization is prone to errors, especially due to environmental changes and illumination variations. To provide a reasonable basis for comparing 5G/mmWave sensing and LiDAR perception capabilities, the authors of [47] proposed a cellular SLAM framework that leverages 5G NR signals and inertial measurements for mobile robot localization using multiple BSs. Their method estimates the robot’s pose while generating a radio signal map based on RSSI measurements for correction purposes. In the study, this approach was benchmarked against LiDAR-inertial odometry smoothing and mapping by comparing its performance against ground-truth references in simulations. Two experimental setups were evaluated using sub-6 GHz and mmWave frequency bands for communication, with transmission based on downlink signals. The results showed that 5G positioning can be effectively used for radio SLAM, improving robustness in outdoor environments and serving as an additional absolute reference source when LiDAR-based methods fail.

### v2

@@ III-C5 Sources of Metric Motion Input
Built upon algorithmic pipelines with the same core elements as in the framework described above and also illustrated in Fig. 6, the perception module of a networked autonomous system may incorporate and augment the vision-based inference mechanisms that have shaped robotic SLAM. Networked implementations reinterpret them within a distributed framework, where visual cues are fused with motion inputs, radio-based signals, and network state information to maintain situational awareness under dynamic conditions.
Figure 6: Abstract representation of V-SLAM as a system subject to the apparent motion dynamics $\mathbf{\dot{X}}_{i}(t)$, fusing image features with motion input from inertial and/or RF-based odometry, and providing landmark and pose estimates.
@@ III-D Active SLAM for Networked Autonomous Systems
III-D1 Communications-driven Active Perception
To provide a unified mathematical framework where a wireless communication objective is viewed as an uncertainty reduction objective and tackled through selecting motion inputs that improve landmark estimation accuracy, the problem of active V-SLAM for a vehicle equipped with a monocular camera operating in an urban environment is considered.
Let $\mathbf{X}_{i,k}$ denote the coordinates of the $i$-th tracked 3-D landmark expressed in the camera coordinate frame at time $k$. In the communication-aware active perception case, the landmark may correspond to a tracked user equipment (UE) or transmitter whose position estimate is important for optimal beam alignment. The apparent landmark dynamics induced by camera motion
$\mathbf{X}_{i,k}\approx\left(\mathbf{I}_{3}+\mathbf{\Omega}_{k}dt\right)\mathbf{X}_{i,k-1}+\mathbf{v}_{k}dt,$ (20)
can be written as:
$\mathbf{X}_{i,k}=\mathbf{f}(\mathbf{X}_{i,k-1},\mathbf{u}_{k})+\mathbf{w}_{i,k},$ (21)
where $\mathbf{u}_{k}$ denotes the velocity/control input $\mathbf{v}_{k}$ over the interval [k1,k][k-1,k], and $\mathbf{w}_{i,k}$ is a term that accounts for process noise.
The feature tracking measurement model for a single visual feature is expressed as
$\mathbf{y}_{i,k}=\mathbf{h}(\mathbf{X}_{i,k})+\mathbf{n}_{i,k},$ (22)
where $\mathbf{h}(\cdot)$ is the camera projection model and $\mathbf{n}_{i,k}$ denotes measurement noise. The expressions $\mathbf{f}(\cdot)$ and $\mathbf{h}(\cdot)$ are consistent with recursive Bayesian filtering, such as the EKF, which maintains the posterior probability of the state estimate
$p_{k}=p(\mathbf{x}_{k}\mid\mathbf{y}_{1:k}),$ (23)
where $\mathbf{x}_{k}$ in this case corresponds to the relative 3-D position of a tracked landmark feature $\mathbf{X}_{i,k}$ and the perspective projection transformation of the feature to the image plane $\mathbf{y}_{i,k}$ is written more compactly as $\mathbf{y}_{k}$.
III-D2 Optimal Control Formulation for Active SLAM
Active SLAM computes a control input that reduces the expected uncertainty of the future estimate. A general optimal control formulation considering a finite-horizon from step 0 to $k_{f}$ is
$\mathbf{u}_{0:k_{f}}^{\star}=\arg\min_{\mathbf{u}_{0:k_{f}}}\mathbb{E}_{\mathbf{y}_{1:k_{f}}}\left[J(p_{k_{f}})\right],$ (24)
where $J(\cdot)$ is the cost function. The posterior probability of the state estimate evolves according to the recursive Bayesian filtering model. The posterior covariance matrix $\mathbf{\Sigma}_{k}$ can be considered for the formalization of an estimation uncertainty metric to minimize such as [23]:
$J(p_{k})=\operatorname{tr}(\mathbf{\Sigma}_{k}),$ (25)
Solving this active SLAM problem online can be computationally demanding. Hence, a one-step approximation can be adopted. Let
$\hat{\mathbf{x}}_{k-1}=\mathbb{E}[\mathbf{x}_{k-1}\mid\mathbf{y}_{1:k-1}]$ (26)
denote the current state estimate, and let
$\hat{\mathbf{x}}_{k}^{-}=\mathbf{f}(\hat{\mathbf{x}}_{k-1},\mathbf{u}_{k})$ (27)
denote the predicted current state before incorporating the new measurement. A gradient-based law approximates $\mathbf{u}_{k}$ so that the state estimate can move along the direction of maximum uncertainty reduction:
$\hat{\mathbf{x}}_{k}^{-}=\hat{\mathbf{x}}_{k-1}-\Gamma\nabla_{\hat{\mathbf{x}}_{k-1}}\mathbb{E}_{\mathbf{y}_{k}}\left[J(p_{k})\mid\mathbf{y}_{1:k-1}\right],$ (28)
where $\Gamma$ is a control gain matrix. For posterior covariance minimization, this becomes [23]:
$\hat{\mathbf{x}}_{k}^{-}=\hat{\mathbf{x}}_{k-1}-\Gamma\nabla_{\hat{\mathbf{x}}_{k-1}}\operatorname{tr}\left(\mathbf{\Sigma}_{k}\right),$ (29)
or equivalently
$\mathbf{f}(\hat{\mathbf{x}}_{k-1},\mathbf{u}_{k})=\hat{\mathbf{x}}_{k-1}-\Gamma\nabla_{\hat{\mathbf{x}}_{k-1}}\operatorname{tr}\left(\mathbf{\Sigma}_{k}\right).$ (30)
By equating the desired evolution of the state estimates with the system dynamics and solving for the control input, the corresponding velocity command at step $k$ can be obtained as
$\mathbf{v}_{k}=\frac{1}{dt}\left(\left(\mathbf{I}_{3}-e^{\mathbf{\Omega}_{k}dt}\right)\hat{\mathbf{X}}_{i,k-1}-\Gamma\nabla_{\hat{\mathbf{X}}_{i,k-1}}\operatorname{tr}\left(\mathbf{\Sigma}_{k}\right)\right).$ (31)
III-D3 Estimation Uncertainty Reduction for Optimal Beam Alignment
The tracked UE or transmitter can be treated as a region of interest (ROI) in the image plane, retrieved using an object detection model [90]. Within this ROI, the feature of interest can be tracked via the LK method. Let $\mathbf{X}_{\mathrm{tx},k}=[X_{\mathrm{tx},k},Y_{\mathrm{tx},k},Z_{\mathrm{tx},k}]^{T}$ denote its estimated 3-D position in the camera coordinate frame. The corresponding LOS path direction is:
$\hat{\mathbf{d}}_{k}=\frac{\mathbf{X}_{\mathrm{tx},k}}{\left\|\mathbf{X}_{\mathrm{tx},k}\right\|_{2}}.$ (32)
The azimuth and elevation angles can then be obtained as
$\phi_{k}=\operatorname{atan2}(Y_{\mathrm{tx},k},X_{\mathrm{tx},k}),$ (33)
$\theta_{k}=\operatorname{atan2}\left(Z_{\mathrm{tx},k},\sqrt{X_{\mathrm{tx},k}^{2}+Y_{\mathrm{tx},k}^{2}}\right).$ (34)
Considering a transmitter / UE with a single omnidirectional antenna and a receiver beam-steering codebook $\mathcal{F}=\{\mathbf{f}_{q}\}_{q=1}^{Q}$, the downlink received signal over the subcarrier $m$ can be modeled as [90]:
$z_{m,k}=\mathbf{h}_{m,k}(\phi_{k},\theta_{k})^{T}\mathbf{f}_{q,k}s_{k}+r_{m,k},$ (35)
where $\mathbf{h}_{m,k}(\phi_{k},\theta_{k})$ is the channel vector, $\mathbf{f}_{q,k}$ is the selected beamforming vector, $s_{k}$ is the transmitted symbol, and $r_{m,k}$ is receiver noise. The optimal beam that aligns better with $\hat{\mathbf{d}}_{k}$ yielding the strongest effective channel gain [90], can be selected as
$q_{k}^{\star}=\arg\max_{q}\left|\mathbf{h}_{m,k}(\phi_{k},\theta_{k})^{T}\mathbf{f}_{q,k}\right|^{2}.$ (36)
Fig. 7 visualizes the active SLAM framework described above.
Figure 7: V-SLAM as an active framework that computes the next motion to achieve the communications objective of beam alignment through estimation uncertainty reduction $\nabla_{\hat{\mathbf{X}}_{\mathrm{tx},k-1}}\operatorname{tr}\left(\mathbf{\Sigma}_{k}\right)$ in the transmitter position $\hat{\mathbf{X}}_{\mathrm{tx},k}$.
III-D4 Limitations of the Active SLAM Uncertainty Reduction-driven Motion
@@
The above formulations demonstrate the capabilities of SLAM as an active estimation framework. Instead of estimating pose and landmarks from available vision and wireless communication measurements, the autonomous vehicle can perform auxiliary motions that could further reduce landmark uncertainty and assist wireless links over LOS paths. For aerial robots, $\mathbf{u}_{k}$ may directly represent a 3-D body-frame velocity command. For ground vehicles / cars, the same communications-driver active perception objective can be applied, however, $\mathbf{u}_{k}$ is restricted through nonholonomic constraints, longitudinal and lateral tire-slip dynamics, and constrained by lane, collision-avoidance, comfort and safety driving decisions. In addition to this, having stopping criteria is imperative, because it can be shown that after reaching a particular point, more information would lead not only to a diminishing return effect but also, in case of ambiguous or contradictory information, to an unrecoverable state (e.g., numerous incorrect loop closures). Lastly, another critical research direction concerns the establishment of mathematical guarantees for active SLAM and the derivation of near-optimal control policies. Given that obtaining an exact, closed form solution to the problem is in the general case computationally heavy, developing approximation algorithms with well-defined performance bounds becomes essential. A representative example of this line of work is the application of submodularity theory [20] in the related domain of active sensor placement, which enables provable near-optimal solutions under certain conditions.
@@ IV-A Mathematical and Algorithmic Methods
Mathematical modeling plays an important role in bridging SLAM with wireless communications. In conventional SLAM, probabilistic estimation techniques help robots and autonomous vehicles navigate and reconstruct their environment. However, when applied to wireless environments, SLAM must account for additional challenges posed by signal propagation, multipath effects, and dynamic radio environments. This interaction between localization, mapping, and wireless signal processing motivates mathematical frameworks that incorporate elements of RF-based sensor fusion, probabilistic inference, signal estimation, and information vectorization.
@@ IV-A1 Probabilistic Approaches in SLAM
Extending previous studies on 5G positioning and environment mapping, [8] explored localization techniques suitable for new radio (NR) and sub-THz 5G communications. They proposed an L1-regularized least-squares approach with low computational complexity to obtain sparse range-angle charts at individual sensing locations. In addition, a novel state model was introduced for tracking diffuse and specular scatterers using an interacting multiple model extended Kalman filter (IMM-EKF) and a smoother. The IMM design included a finite set of motion models with probabilistic transitions between them, following the principles of mmWave SLAM. Experimental results demonstrated the effectiveness of the proposed methods, with IMM-EKF-based tracking significantly outperforming static modeling approaches.
@@ IV-A3 Vector-field SLAM
A vector field is defined as a regular grid with fixed node positions $\mathbf{b}_{i}$, where $i=\{1,\dots,N\}$, while each node $\mathbf{m}_{i}\in\mathbb{R}^{M}$ represents a vector of expected signal values when the robot is positioned at $\mathbf{b}_{i}$ with a fixed orientation $\theta_{0}$. Without loss of generality, let $\theta_{0}=0$. The vector field and robot pose are then estimated as follows.
@@
$\mathbf{x}_{t}=\begin{bmatrix}x(t)\\ y(t)\\ \theta(t)\end{bmatrix}$. (37)
@@
$\mathbf{x}_{0}=\begin{bmatrix}0\\ 0\\ 0\end{bmatrix}$. (38)
At each time step $t=\{1,\dots,T\}$, the robot executes a motion/control input $\mathbf{u}_{t}$ with covariance $\mathbf{R}_{t}$ and receives a measurement $\mathbf{z}_{t}$ of the continuous robot state signals with covariance $\mathbf{Q}_{t}$.
@@
$\mathbf{x}_{t}=\mathbf{g}(\mathbf{x}_{t-1},\mathbf{u}_{t})+\mathbf{e}_{u},$ (39)
@@
$\mathbf{z}_{t}=\mathbf{h}\left(\mathbf{x}_{t},\begin{bmatrix}\mathbf{m}_{1},&\dots,&\mathbf{m}_{N}\end{bmatrix}^{T}\right)+\mathbf{e}_{z},$ (40)
@@
$h\!\left(\mathbf{x}_{t},[\mathbf{m}_{1},\ldots,\mathbf{m}_{N}]^{T}\right)=h_{R}\!\left(h_{0}\!\left(x,y,[\mathbf{m}_{1},\ldots,\mathbf{m}_{N}]^{T}\right),\theta\right),$ (41)
@@
$h_{0}\left(x,y,\begin{bmatrix}\mathbf{m}_{1},&\dots,&\mathbf{m}_{N}\end{bmatrix}^{T}\right)=\sum_{j=0}^{3}w_{j}\mathbf{m}_{i_{j}}.$ (42)
@@
Table III summarizes the key characteristics of the discussed mathematical and algorithmic methods.
TABLE III: Comparison of Representative Mathematical and Algorithmic SLAM Methods
SLAM Method Category Key Characteristics Application Domain / Signal Modalities Advantages Constraints Representative Works
Probabilistic Approaches in SLAM Likelihood-based modeling accounting for multipath effects (specular/diffuse); multi-model estimation (filter banks, particle filtering); recursive Bayesian inference. 5G/mmWave positioning; RFID-based localization; inertial & RSS fusion. High accuracy and robustness; relaxation of Gaussianity assumption; robust under limited prior information. High computational complexity; particle degeneracy; scalability challenges in dense environments. [35, 8, 37, 134, 51, 124]
Graph-based SLAM Factor graph formulation for global optimization; pose graph construction from motion constraints; trajectory refinement via optimization. Visible light positioning; monocular/stereo SLAM; RSS-based localization; fusion with inertial data. Improves global map consistency; reduces drift via loop closure; supports large-scale mapping. Computational overhead in large graphs; sensitivity to incorrect loop closure. [141]
Vector Field SLAM Continuous spatial signal field representation; RF signals modeled as vector-valued fields; interpolation-based observation models. WiFi RSS-based positioning; RF-based SLAM; structured light pattern-based positioning. Suitable for signal-based mapping; no explicit landmark extraction required. Assumes signal stationarity; requires identifiable signals; sensitive to dynamic environments. [39]
Lessons Learned:
  1. 1.
    Modeling Fidelity: Likelihood-based formulations can enable the incorporation of the complex scattering nature of RF propagation, yielding high accuracy and robustness. Factor graph representations can naturally encode geometric and motion constraints in visual and inertial systems, resulting in enhanced global map consistency and reduced drift in pose estimation. Vector field formulations enable high-fidelity signal-driven modeling without the need for explicit determination of landmarks.
  2. 2.
    Complexity: Probabilistic and graph-based approaches can incur significant computational cost in large-scale, dense environments. Vector field methods alleviate the need for explicit modeling through functional approximations, at the expense of stronger assumptions made on signal stationarity and identifiability.
IV-B Sensor Modality-Driven SLAM Techniques
@@ IV-B2 LiDAR-based SLAM
In [12], the authors investigated hardware and software co-design strategies to meet the requirements of mobile robot localization and object perception. To enable accurate mobile positioning, they proposed a SLAM strategy that integrates LiDAR, a monocular camera, and an IMU. For object perception, the study exemplifies the extraction of meter information using DL techniques for meter detection, localization, and text recognition. The integration of 5G and WiFi communication enables an intelligent robotic inspection system with a network management platform. The authors of [131] addressed the challenge of robust long-term localization in cluttered indoor environments, such as offices and corridors, using LiDAR point clouds within an area graph. This hierarchical, topometric semantic map representation uses polygons to define spaces such as rooms, corridors, or buildings. Their approach, called area graph localization (AGloc), is augmented by a clutter removal subsampling method to remove excessive noise, as well as sampling techniques and weighting functions for global robot localization (Fig. 10). First, a set of candidate robot positions is generated using a sampling technique, after which a scoring function evaluates each candidate based on its match with the area graph. The point-to-point iterative closest point (ICP) algorithm and a weighting function enable localization using only wall and passage representations, ensuring robust lifelong localization despite the presence of non-static obstacles such as furniture, pets, and humans. Additionally, a ’corridorness’ score function is introduced to provide downsampled point clouds, improving ICP localization in corridor-like environments.
@@
In [155], a high-precision multimodal approach for indoor positioning was proposed using LiDAR as the central sensor modality. This method improves localization accuracy by incorporating widely available technologies, such as WiFi access points (APs) and low-cost cameras. In this hybrid scheme, positioning data from LiDAR-based SLAM is used as a theoretical reference for visual images, while WiFi signals estimate the rough location through likelihood. Finally, RGB images refine the ground map and improve the localization accuracy. As mentioned above, LiDAR-based localization is prone to errors, especially due to environmental changes and illumination variations. To provide a reasonable basis for comparing 5G/mmWave sensing and LiDAR perception capabilities, the authors of [49] proposed a cellular SLAM framework that leverages 5G NR signals and inertial measurements for mobile robot localization using multiple BSs. Their method estimates the robot’s pose while generating a radio signal map based on RSSI measurements for correction purposes. In the study, this approach was benchmarked against LiDAR-inertial odometry smoothing and mapping by comparing its performance against ground-truth references in simulations. Two experimental setups were evaluated using sub-6 GHz and mmWave frequency bands for communication, with transmission based on downlink signals. The results showed that 5G positioning can be effectively used for radio SLAM, improving robustness in outdoor environments and serving as an additional absolute reference source when LiDAR-based methods fail.